# Cross-modal RAG: Sub-dimensional Retrieval-Augmented Text-to-Image Generation

### v2

@@ Abstract
Text-to-image generation increasingly demands access to domain-specific, fine-grained, and rapidly evolving knowledge that pretrained models cannot fully capture. Existing Retrieval-Augmented Generation (RAG) methods attempt to address this by retrieving globally relevant images, but they fail when no single image contains all desired elements from a complex user query. We propose Cross-modal RAG, a novel framework that decomposes both queries and images into sub-dimensional components, enabling subquery-aware retrieval and generation. Our method introduces a hybrid retrieval strategy—combining a sub-dimensional sparse retriever with a dense retriever—to identify a Pareto-optimal set of images, each contributing complementary aspects of the query. During generation, a multimodal large language model is guided to selectively condition on relevant visual features aligned to specific subqueries, ensuring subquery-aware image synthesis. Extensive experiments on MS-COCO, Flickr30K, WikiArt, CUB, and ImageNet-LT demonstrate that Cross-modal RAG significantly outperforms existing baselines in both retrieval and generation quality, while maintaining high efficiency.
@@ 1 Introduction
Text-to-Image Generation (T2I-G) has witnessed rapid progress in recent years, driven by advances in diffusion models rombach2022high ; saharia2022photorealistic ; nichol2021glide and multimodal large language models (MLLMs) openai2025gpt4oimage ; google2025gemini2flash ; jin2023unified ; zhu2024latentexplainer ; bai2024beyond , enabling the synthesis of increasingly realistic and diverse images from natural language descriptions. However, in many real-world applications, domain-specific image generation requires knowledge that is not readily encoded within pre-trained image generators, especially when such information is highly long-tailed, fast-updated, and proprietary. To address this limitation, Retrieval-Augmented Generation (RAG) has emerged as a promising paradigm by incorporating an exrternal image database to supply factual reference during generation zheng2025retrieval ; zhao2024retrieval . Notable RAG-based image generation approaches such as Re-Imagen chen2022re , RDM blattmann2022retrieval , and KNN-Diffusion sheynin2022knn integrate retrieved images with diffusion models to improve output fidelity. However, these existing RAG methods typically rely on off-the-shelf retrievers (e.g., those based on CLIP radford2021learning ) which compute global image-text similarities and retrieve whole images based on the full user query. This coarse-grained retrieval strategy often fails in complex scenarios where the query involves multiple fine-grained entities or attributes varma2023villa – especially when no single image contains all required components in the query. In practice, it is extremely common that single images in the retrieval database only satisfy a subset of the query. As in Fig. 1, no single image in the retrieval database perfectly covers all four aspects in the query; instead, each covers different subsets of the query. Existing RAG methods often retrieve top-k images based on the entire query, so images that redundantly contain most aspects of the query tend to be retrieved (e.g., three images that all include “Cybertruck” and “Tesla store”), while some aspects may be underweighted (e.g., “Venus ground”) or even missed (e.g., “third-generation Labubu”), leading to the distortion in the missed aspects. Also, during generation, existing RAG has not be precisely instructed about which aspects of each image should be leveraged, resulting in the superfluous lightning in the generated image by previous RAG.
@@
Therefore, instead of being restricted to considering whether each whole image is related to the entirety of the query, it is desired to advance to pinpointing which aspects (i.e., sub-dimensions) of which images can address which aspects of the query for image generation. Such desire is boiled down to several open questions. First, how to precisely gauge which part of the queries match which aspects of each image? Existing global embedding methods, like CLIP, do not naturally support sub-dimensional alignment radford2021learning , and current fine-grained vision-language matching is limited to region-level object patterns varma2023villa ; zhong2022regionclip , which are computationally expensive and error-prone. Second, how to retrieve the smallest number of images that cover all necessary information? It is desired to retrieve an optimal set of images such that each covers different aspects of the query while avoiding redundancy in order to maximize the amount of relevant information under a limited context window size. Third, how to precisely inform the image generator of what aspect of each image to refer to when generating? Existing image generators typically can take in the input query or images, but here it requires adding fine-grained instructions about how to leverage relevant aspects of images when generating, which is not well explored.
To address these open problems, we propose Cross-modal Sub-dimensional Retrieval Augmented Generation (Cross-modal RAG), a novel text-to-image generation that can identify, retrieve, and leverage image sub-dimensions to satisfy different query aspects. To decompose and identify key image sub-dimensions, we decompose the user query into subqueries and candidate images into sub-dimensional representations with respect to the subqueries, enabling accurate subquery-level alignment. To retrieve comprehensive and complementary image sub-dimensions, we formulate the retrieval goal as a multi-objective optimization problem and introduce an efficient hybrid retrieval strategy – combining a lightweight sub-dimensional sparse retriever with a sub-dimensional dense retriever – to retrieve a set of Pareto-optimal images that collectively cover all subqueries in the query as in Fig. 1 (right). To effectively instruct image generators with the retrieved image sub-dimensions, we present a model-agnostic and subquery-aware generation with MLLMs, which explicitly preserves and composes the subquery-aligned components from the retrieved images into a coherent final image. For instance, our method only preserves the “Venus ground” in the final image, while previous RAG can also preserve the irrelevant lightning in Fig. 1. Extensive experiments demonstrate that Cross-modal RAG achieves state-of-the-art performance in both text-to-image retrieval and text-to-image generation tasks across multiple fine-grained, domain-specific, and long-tailed image benchmarks, while maintaining excellent computational efficiency.
@@ 2.1 Text-to-Image Generation
Text-to-Image Generation (T2I-G) has made significant strides, evolving through methodologies such as Generative Adversarial Networks (GANs) goodfellow2014generative ; brock2018large , auto-regressive models van2016pixel ; ramesh2021zero , and diffusion models ho2020denoising ; nichol2021improved . Recent breakthroughs in diffusion models and multimodal large language models (MLLMs), driven by scaling laws kaplan2020scaling , have significantly advanced the capabilities of T2I-G. Notable examples include the DALL-E series ramesh2021zero ; betker2023improving , the Imagen series saharia2022photorealistic , and the Stable Diffusion (SD) series rombach2022high ; podell2023sdxl ; esser2024scaling . More recently, image generation functionalities have been integrated directly into advanced MLLMs such as GPT Image openai2025gpt4oimage and Gemini 2.0 Flash Image Generation google2025gemini2flash . However, despite these advancements, traditional T2I-G methods often struggle with knowledge-intensive, long-tailed, and fine-grained image-generation tasks. These scenarios typically require additional context to generate accurate images, necessitating RAG techniques.
@@ 2.2 Text-to-Image Retrieval
Text-to-Image Retrieval (T2I-R) has become a crucial subtask in supporting fine-grained image understanding and generation. CLIP radford2021learning is currently the most widely adopted approach, mapping images and texts into a shared embedding space via contrastive learning. While CLIP excels at coarse-grained alignment, it underperforms in fine-grained text-to-image retrieval, especially in scenes involving multiple objects or nuanced attributes. ViLLA varma2023villa explicitly highlights this limitation, demonstrating that CLIP fails to capture detailed correspondences between image regions and textual attributes. SigLIP zhai2023sigmoid , along with other refinements such as FILIP yao2021filip and SLIP mu2022slip , improves CLIP’s contrastive learning framework and achieves superior zero-shot classification performance. However, these methods still rely on global image-text embeddings, which are inadequate for resolving localized visual details required by fine-grained queries.
To address this, recent works on fine-grained text-to-image retrieval (e.g., ViLLA varma2023villa , RegionCLIP zhong2022regionclip ) have adopted region-based approaches that involve cropping image patches for localized alignment. In contrast, our vision-based sub-dimensional dense retriever bypasses the need for explicit cropping. By constructing sub-dimensional vision embeddings directly from the full image, we enable more efficient and effective matching against subqueries.
@@ 2.3 Retrieval-Augmented Generation
Retrieval-Augmented Generation has demonstrated significant progress in improving factuality for both natural language generation gao2023retrieval ; zhu2023large and image generation chen2022re ; yasunaga2022retrieval . Most RAG-based approaches for image generation are built upon diffusion models (e.g., Re-Imagen chen2022re , RDM blattmann2022retrieval , KNN-Diffusion sheynin2022knn ), but these methods largely overlook fine-grained semantic alignment. FineRAG yuan2025finerag takes a step toward fine-grained image generation by decomposing the textual input into fine-grained entities; however, it does not incorporate fine-grained decomposition on the visual side. In contrast, our approach performs dual decomposition: (i) the query is parsed into subqueries that capture distinct semantic components, and (ii) the candidate images are decomposed into sub-dimensional vision embeddings aligned with the corresponding subqueries. Furthermore, while existing RAG-based image models typically rely on off-the-shelf retrievers, we introduce a novel retrieval method that combines a sub-dimensional sparse filtering stage with a sub-dimensional dense retriever. Finally, with the recent surge of MLLM-based image generation, we explore how our fine-grained retrieval information can be integrated to guide generation at the sub-dimensional level.
@@ 3.1.1 Sub-dimensional Dense Retriever
Here, the similarity score $sim$ is cosine similarity. The similarity scores are aggregated across $n$ subqueries to form an overall similarity metric $S(Q,I_{j})$. Images are ranked based on their similarity $S(Q,I_{j})$ for the given query, and the top-ranked images are retrieved.
$v_{ji}=f_{a}\left(\Phi_{\text{clip-v}}(I_{j}),t_{i}\right).$ (2)
The vision adapter consists of: 1) A multi-head vision cross-attention layer where the learnable query tokens attend to the vision embeddings extracted from the frozen CLIP visual encoder; 2) A multi-head text cross-attention layer where the output of the vision cross-attention further attends to the subquery embeddings extracted from the frozen CLIP text encoder; 3) An MLP head that maps the attended features to a shared multimodal embedding space, followed by layer normalization. The output $v_{ji}$ represents $I_{j}$’s ith-dimensional vision embedding corresponding to the subquery $q_{i}$, which is decomposed from $Q$ and can be obtained by an off-the-shelf LLM (e.g. GPT-4o mini) using the structured prompt in Appendix A.
@@ 3.1.2 Sub-dimensional Sparse Retriever
$s_{i}\left(I_{j}\right)=\begin{cases}1,&\text{ if the caption of }I_{j}\text{ contains }q_{i}\\ 0,&\text{ otherwise }\end{cases}$ (5)
@@ 3.1.3 Multi-objective Optimization Formulation and Algorithm
Each subquery can be regarded as a distinct objective, giving rise to a multi-objective optimization problem for retrieval. Our primary goal is to select images that collectively maximize text-based subquery satisfaction (sub-dimensional sparse retrieval), while also maximizing fine-grained vision-based similarity (sub-dimensional dense retrieval). Thus, the overall objective is formalized as:
@@
1:Query $Q$ decomposed into subqueries $\{q_{i}\}_{i=1}^{n}$, image retrieval database $\mathcal{D}$, weights $\{\alpha_{i}\}_{i=1}^{n}$ with $\alpha_{i}>0,\sum_{i}\alpha_{i}=1$, trade-off parameter $\beta$ with $0<\beta<\beta_{max}$
@@
6:$\widetilde{D}\leftarrow\{I_{j}\mid\mathbf{s}\left(I_{j}\right)\text{ is not all zeros}\}$
7:$\mathcal{P}\leftarrow\emptyset$
@@ 3.2 Image Generation with Retrieved Images
To generate an image from the user query $Q$ while ensuring that the satisfied subquery features in $\mathcal{R}$ are preserved, we utilize a pretrained MLLM with subquery-aware instructions.
@@
For each image $I_{j}^{*}\in\mathcal{P}$, we construct an in-context example in a form: $\langle I_{j}^{*}\rangle$ Use only [$Q_{r}$] in [$I_{j}^{*}$].
Here, $\langle I_{j}^{*}\rangle$ denotes the visual tokens for the $r$-th retrieved image, [$Q_{r}$] is the satisfied subqueries in $I_{j}^{*}$, and [$I_{j}^{*}$] is "the $r$-th retrieved image".
Next, we feed the in-context examples to a pretrained MLLM together with the original query $Q$. The MLLM, which operates in an autoregressive manner, is thus guided to generate the final image $\hat{I}$ as:
$p_{\theta}\bigl{(}\hat{I}\;\big{|}\;Q,\{I_{j}^{*}\},\{Q_{r}\}\bigr{)}\;=\;\prod_{t=1}^{T}p_{\theta}\Bigl{(}\hat{I}_{t}\,\Big{|}\,\hat{I}_{<t},\,Q,\,\{I_{j}^{*}\},\{Q_{r}\}\bigr{)},$ (10)
where $\hat{I}_{t}$ denotes the $t$-th visual token in the generated image representation, and $\theta$ represents the parameters of the pretrained MLLM. By referencing the full prompt: [$Q$] $\langle I_{j}^{*}\rangle\;\texttt{Use only [}Q_{r}\texttt{] in [}I_{j}^{*}\texttt{]}$, the MLLM learns to preserve the relevant subquery features that each retrieved image $I_{j}^{*}$ contributes.

### v3

# Cross-modal RAG: Sub-dimensional Text-to-Image Retrieval-Augmented Generation

@@ Abstract
Text-to-image generation increasingly demands access to domain-specific, fine-grained, and rapidly evolving knowledge that pretrained models cannot fully capture, necessitating the integration of retrieval methods. Existing Retrieval-Augmented Generation (RAG) methods attempt to address this by retrieving globally relevant images, but they fail when no single image contains all desired elements from a complex user query. We propose Cross-modal RAG, a novel framework that decomposes both queries and images into sub-dimensional components, enabling subquery-aware retrieval and generation. Our method introduces a hybrid retrieval strategy—combining a sub-dimensional sparse retriever with a dense retriever—to identify a Pareto-optimal set of images, each contributing complementary aspects of the query. During generation, a multimodal large language model is guided to selectively condition on relevant visual features aligned to specific subqueries, ensuring subquery-aware image synthesis. Extensive experiments on MS-COCO, Flickr30K, WikiArt, CUB, and ImageNet-LT demonstrate that Cross-modal RAG significantly outperforms existing baselines in the retrieval and further contributes to generation quality, while maintaining high efficiency.
@@ 1 Introduction
Text-to-Image Generation (T2I-G) has witnessed rapid progress in recent years, driven by advances in diffusion models (Rombach et al., 2022; Saharia et al., 2022; Nichol et al., 2021) and multimodal large language models (MLLMs) (OpenAI, 2025; Google, 2025; Jin et al., 2023), enabling the synthesis of increasingly realistic and diverse images from natural language descriptions. However, in many real-world applications, domain-specific image generation requires knowledge that is not readily encoded within pre-trained image generators, especially when such information is highly long-tailed, fast-updated, and proprietary. To address this limitation, Retrieval-Augmented Generation (RAG) has emerged as a promising paradigm by incorporating an external image database to supply factual reference during generation (Zheng et al., 2025; Zhao et al., 2024). Notable RAG-based image generation approaches such as Re-Imagen (Chen et al., 2022), RDM (Blattmann et al., 2022), and KNN-Diffusion (Sheynin et al., 2022) integrate retrieved images with diffusion models to improve output fidelity. However, these existing RAG methods typically rely on off-the-shelf retrievers (e.g., those based on CLIP (Radford et al., 2021)) which compute global image-text similarities and retrieve whole images based on the full user query. This coarse-grained retrieval strategy often fails in complex scenarios where the query involves multiple fine-grained entities or attributes (Varma et al., 2023) – especially when no single image contains all required components in the query. In practice, it is extremely common that single images in the retrieval database only satisfy a subset of the query. As in Fig. 1, no single image in the retrieval database perfectly covers all four aspects in the query; instead, each covers different subsets of the query. Existing RAG methods often retrieve top-k images based on the entire query, so images that redundantly contain most aspects of the query tend to be retrieved (e.g., three images that all include “Cybertruck” and “Tesla store”), while some aspects may be underweighted (e.g., “Venus ground”) or even missed (e.g., “third-generation Labubu”), leading to the distortion in the missed aspects. Also, during generation, existing RAG has not been precisely instructed about which aspects of each image should be leveraged, resulting in the superfluous lightning in the generated image by previous RAG.
@@
Therefore, instead of being restricted to considering whether each whole image is related to the entirety of the query, it is desired to advance to pinpointing which aspects (i.e., sub-dimensions) of which images can address which aspects of the query for image generation. Such desire is boiled down to several open questions. First, how to precisely gauge which part of the queries match which aspects of each image? Existing global embedding methods, like CLIP, do not naturally support sub-dimensional alignment (Radford et al., 2021), and current fine-grained vision-language matching is limited to region-level object patterns (Varma et al., 2023; Zhong et al., 2022), which are computationally expensive and error-prone. Second, how to retrieve the smallest number of images that cover all necessary information? It is desired to retrieve an optimal set of images such that each covers different aspects of the query while avoiding redundancy in order to maximize the amount of relevant information under a limited context window size. Third, how to precisely inform the image generator of what aspect of each image to refer to when generating? Existing image generators typically can take in the input query or images, but here it requires adding fine-grained instructions about how to leverage relevant aspects of images when generating, which is not well explored.
To address these open problems, we propose Cross-modal Sub-dimensional Retrieval Augmented Generation (Cross-modal RAG), a novel text-to-image retrieval-augmented framework that can identify, retrieve, and leverage image sub-dimensions to satisfy different query aspects:
  • To decompose and identify key image sub-dimensions, we decompose the user query into subqueries and candidate images into sub-dimensional representations with respect to the subqueries, enabling accurate subquery-level alignment.
  • To retrieve comprehensive and complementary image sub-dimensions, we formulate the retrieval goal as a multi-objective optimization problem and introduce an efficient hybrid retrieval strategy – combining a lightweight sub-dimensional sparse retriever with a sub-dimensional dense retriever – to retrieve a set of Pareto-optimal images that collectively cover all subqueries in the query as in Fig. 1 (right).
  • To effectively instruct image generators with the retrieved image sub-dimensions, we present a model-agnostic and subquery-aware generation with MLLMs, which explicitly preserves and composes the subquery-aligned components from the retrieved images into a coherent final image. For instance, our method only preserves the “Venus ground” in the final image, while previous RAG can also preserve the irrelevant lightning in Fig. 1.
Extensive experiments demonstrate that Cross-modal RAG achieves state-of-the-art performance in the text-to-image retrieval and also benefits generation tasks across multiple fine-grained, domain-specific, and long-tailed image benchmarks, while maintaining excellent computational efficiency.
@@ 2.1 Text-to-Image Generation
Text-to-Image Generation (T2I-G) has made significant strides, evolving through methodologies such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Brock et al., 2018), auto-regressive models (Van Den Oord et al., 2016; Ramesh et al., 2021), and diffusion models (Ho et al., 2020; Nichol & Dhariwal, 2021). Recent breakthroughs in diffusion models and multimodal large language models (MLLMs), driven by scaling laws (Kaplan et al., 2020), have significantly advanced the capabilities of T2I-G. Notable examples include the DALL-E series (Ramesh et al., 2021; Betker et al., 2023), the Imagen series (Saharia et al., 2022), and the Stable Diffusion (SD) series (Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024). More recently, image generation functionalities have been integrated directly into advanced MLLMs such as GPT Image (OpenAI, 2025) and Gemini Image Generation (Google, 2025). However, despite these advancements, traditional T2I-G methods often struggle with knowledge-intensive, long-tailed, and fine-grained image-generation tasks. These scenarios typically require additional context to generate accurate images, necessitating RAG techniques.
@@ 2.2 Text-to-Image Retrieval
Text-to-Image Retrieval (T2I-R) has become a crucial subtask in supporting fine-grained image understanding and generation. CLIP (Radford et al., 2021) is currently the most widely adopted approach, mapping images and texts into a shared embedding space via contrastive learning. While CLIP excels at coarse-grained alignment, it underperforms in fine-grained text-to-image retrieval, especially in scenes involving multiple objects or nuanced attributes. ViLLA (Varma et al., 2023) explicitly highlights this limitation, demonstrating that CLIP fails to capture detailed correspondences between image regions and textual attributes. SigLIP (Zhai et al., 2023), along with other refinements such as FILIP (Yao et al., 2021) and SLIP (Mu et al., 2022), improves CLIP’s contrastive learning framework and achieves superior zero-shot classification performance. However, these methods still rely on global image-text embeddings, which are inadequate for resolving localized visual details required by fine-grained queries.
To address this, recent works on fine-grained text-to-image retrieval (e.g., ViLLA (Varma et al., 2023), RegionCLIP (Zhong et al., 2022)) have adopted region-based approaches that involve cropping image patches for localized alignment. In contrast, our vision-based sub-dimensional dense retriever bypasses the need for explicit cropping. By constructing sub-dimensional vision embeddings directly from the full image, we enable more efficient and effective matching against subqueries.
@@ 2.3 Retrieval-Augmented Generation
Retrieval-Augmented Generation has demonstrated significant progress in improving factuality for both natural language generation (Gao et al., 2023; Zhu et al., 2023) and image generation (Chen et al., 2022; Yasunaga et al., 2022). Most RAG-based approaches for image generation are built upon diffusion models (e.g., Re-Imagen (Chen et al., 2022), RDM (Blattmann et al., 2022), KNN-Diffusion (Sheynin et al., 2022)), but these methods largely overlook fine-grained semantic alignment. FineRAG (Yuan et al., 2025) takes a step toward fine-grained image generation by decomposing the textual input into fine-grained entities; however, it does not incorporate fine-grained decomposition on the visual side. In contrast, our approach performs dual decomposition: (i) the query is parsed into subqueries that capture distinct semantic components, and (ii) the candidate images are decomposed into sub-dimensional vision embeddings aligned with the corresponding subqueries. Furthermore, while existing RAG-based image models typically rely on off-the-shelf retrievers, we introduce a novel retrieval method that combines a sub-dimensional sparse filtering stage with a sub-dimensional dense retriever. Finally, with the recent surge of MLLM-based image generation, we explore how our fine-grained retrieval information can be integrated to guide generation at the sub-dimensional level.
@@ 3.1.1 Sub-dimensional Dense Retriever
Here, the similarity score simsim is cosine similarity. The similarity scores are aggregated across $n$ subqueries to form an overall similarity metric $S(Q,I_{j})$. Images are ranked based on their similarity $S(Q,I_{j})$ for the given query, and the top-ranked images are retrieved.
@@
$v_{ji}=f_{a}\left(\Phi_{\text{clip-v}}(I_{j}),T_{i}\right).$ (3)
The vision adapter consists of: 1) A multi-head vision cross-attention layer where the learnable query tokens attend to the vision embeddings extracted from the frozen CLIP visual encoder; 2) A multi-head text cross-attention layer where the output of the vision cross-attention further attends to the subquery embeddings extracted from the frozen CLIP text encoder; 3) An MLP head that maps the attended features to a shared multimodal embedding space, followed by layer normalization. The output $v_{ji}$ represents $I_{j}$’s ith-dimensional vision embedding corresponding to the core concept of subquery $q_{i}$, which is decomposed from $Q$ and can be obtained by an off-the-shelf LLM (e.g. GPT-4o mini) using the structured prompt in Appendix A. $T_{i}$ denotes the embedding of the core concept extracted from $q_{i}$, representing the object category without attribute modifiers.
@@ 3.1.2 Sub-dimensional Sparse Retriever
[…]
@@ 3.1.3 Multi-objective Optimization Formulation and Algorithm
Each subquery can be regarded as a distinct objective, giving rise to a multi-objective optimization problem for retrieval. Our primary goal is to select images that collectively maximize text-based subquery satisfaction (sub-dimensional sparse retrieval), while also maximizing fine‐grained vision-based similarity (sub-dimensional dense retrieval). Thus, the overall objective is formalized as:
@@
1:Query $Q$ decomposed into subqueries $\{q_{i}\}_{i=1}^{n}$, image retrieval database $\mathcal{D}$, weights $\{\alpha_{i}\}_{i=1}^{n}$ with $\alpha_{i}>0,\sum_{i}\alpha_{i}=1$, trade‑off parameter $\beta$ with $0<\beta<\beta_{max}$
@@
[…]
@@ 3.2 Image Generation with Retrieved Images
To generate an image from the user query $Q$ while ensuring that the satisfied subquery features in $\mathcal{P}$ are preserved, we utilize a pretrained MLLM with subquery-aware instructions.
@@
For each image $I_{j}^{*}\in\mathcal{P}$, we construct an in‐context example in a form: $\langle I_{j}^{*}\rangle$ Use only $Q_{j}$] in $I_{j}^{*}$].
Here, $\langle I_{j}^{*}\rangle$ denotes the visual tokens for the $j$-th retrieved image, $Q_{j}$] is the satisfied subqueries in $I_{j}^{*}$, and $I_{j}^{*}$] is "the $j$-th retrieved image".
Next, we feed the in‐context examples to a pretrained MLLM together with the original query $Q$. The MLLM, which operates in an autoregressive manner, is thus guided to generate the final image $\hat{I}$ as:
$p_{\theta}\bigl(\hat{I}\;\big|\;Q,\{I_{j}^{*}\},\{Q_{j}\}\bigr)\;=\;\prod_{t=1}^{T}p_{\theta}\Bigl(\hat{I}_{t}\,\Big|\,\hat{I}_{<t},\,Q,\,\{I_{j}^{*}\},\{Q_{j}\}\bigr),$ (10)
[…]